OpenStudy (haleyelizabeth2017):

The design of a digital box camera maximizes the volume while keeping the sum of the dimensions at 6 inches. If the length must be 1.5 times the height, what should each dimension be? *Hint: Let x represent one of the dimensions, and then define the other dimensions in terms of x.

OpenStudy (anonymous):

does the "sum of the dimensions" mean base time width times height? if so then you can write \[x+1.5x+y=6\] or \[2.5x+y=6\] solve that for \(y\)

OpenStudy (haleyelizabeth2017):

I'm just not sure how to write the equations.....something that has to do with 1.5x + something =6....

OpenStudy (haleyelizabeth2017):

okay that would be y=-2.5x+6

OpenStudy (haleyelizabeth2017):

I'm not sure....what I put is all they gave me :(

OpenStudy (anonymous):

then the volume is \[V(x)=x\times 1.5x\times (-2.5x+6)\]

OpenStudy (haleyelizabeth2017):

okay so that is just like what I'm used to! :D

OpenStudy (anonymous):

probably easier to use fractions instead of decimals in any case you can minimize that one using calc

OpenStudy (haleyelizabeth2017):

I like decimals better :P lol idk why........but I am going to do x*1.5x first which is \(1.5x^2\) then I multiply that by -2.5x+6.........which would equal \(-3.75x^3+9x\) correct?

OpenStudy (haleyelizabeth2017):

how does finding that help to find the dimensions?